OpenStudy (anonymous):

Write the product as a trinomial. (x + 3)(x + 4) x2+ 7 x2 + 7x + 12 x2 + 12x + 7 x2 + 12x + 12

OpenStudy (anonymous):

expand it

OpenStudy (anonymous):

x*(x+4)+3*(x+4) x^2+4x+3x+12 =x^2+7x+12

OpenStudy (anonymous):

but where did you get 12 from, that's what i'm confused about

OpenStudy (anonymous):

@g17

OpenStudy (anonymous):

3*(x+4)=3*x+3*4=3x+12